OpenStudy (anonymous):

Describe the differences between the graph of y = -3(x + 7)2 - 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

@elementwielder @terenzreignz @amistre64

OpenStudy (anonymous):

is that x times 2?

OpenStudy (anonymous):

the 1st one has a negative slope 2nd has y-intercept of 0

OpenStudy (anonymous):

\[y=-3(x+7)^2-10\] @elementwielder

OpenStudy (anonymous):

y=-3x^2-42x-157 ok so the equation above is a parabola where as the other one is a regular line

OpenStudy (amistre64):

one is a modified version of the other, your material covers all the modifications and how the affect the parent function

OpenStudy (amistre64):

x^2 is shifted by h: (x-h)^2 it is made skinny or wide by a factor of 'a': a(x-h)^2 it is then lifted or lowered by k: a(x-h)^2 + k

OpenStudy (anonymous):

i was right though... 1st is parabola 2nd is regular?

OpenStudy (anonymous):

(x^2-4x+4)+10(y^2+4+4)=-16+16+40 4(x-2)^2+10(y+2)^2=40 (x-2)^2/10+(y+2)^2/4=40 ... Standard form of ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, (a>b), with (h,k) being ... The difference between the two forms is the interchange of a^2 and b^2.